### 16 Mar 2026 v1

# Rigidity and Structural Asymmetry of Bounded Solutions

Abstract
In this manuscript, we introduce a family of parametrized non-homogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter. We identify a Rotation number hypothesis on the non-homogeneous term, which establishes a structural asymmetry: if two solutions with the same initial condition equal to 11, corresponding respectively to the parameters $s$ and $1-\overline{s}$ lying in the critical strip, are both bounded on $1,+\infty)$, then $\Re(s)=\tfrac{1}{2}$.
Keywords: Non-homogeneous complex linear differential equation, bounded solutions, Volterra integral equation.
MSC subject classifications: 34A30, 34E05
1 Main result
In this manuscript, we study the following non-homogeneous linear complex differential equations:
$\dot{\theta}=wt^{-1}\theta+t^{-1}\eta(t),\quad\theta(1)=\frac{1}{1-w},\quad\theta:[1,+\infty)\to\mathbb{C},$ (1)
where $w\in\mathbb{C}$ such that $\Im(w)\neq 0$ is the parameter and the non-homogeneous term $\eta\in L^{\infty}(\mathbb{C})$ where $L^{\infty}(\mathbb{C})$ is the set of bounded and locally integrable functions from $1,+\infty)$ to $\mathbb{C}$. As formulated in the Dynamical Conjecture in [1], we assume that the function $\eta$ satisfies the following Rotation number hypothesis:
$\exists\rho\in\mathbb{C}^{*}:\quad\sup_{t\geq 1}\Big|\int_{1}^{t}\,\big(\eta(u)-\rho\big)du\Big|<+\infty,$ (H)
The complex number $\rho$ is called the rotation number of $\eta$.
We are interested in the initial conditions of the bounded solutions of the previous differential equation. To this end, we analyze the behavior of the transformed solution $t\mapsto(1-w)t^{-i\Im(w)}\theta(t)$, which leads us to consider the following differential equation:
$\dot{x}=\Re(w)t^{-1}x+(1-w)t^{-1-i\Im(w)}\eta(t),$ (2)
$t\in[1,+\infty),\quad x(1)=1,\quad x:[1,+\infty)\to\mathbb{C}.$
Since $\eta\in L^{\infty}(\mathbb{C})$, the function
$t\mapsto t^{\Re(w)}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1.$
is absolutely continuous on $1,+\infty)$. The differential equation (2) is a non-homogeneous linear differential equation. Then, there exists a unique continuous solution $\psi_{\eta,w}(\cdot)$ of (2) such that $\psi_{w}(1)=1$, which is given by
$\psi_{\eta,w}(t)=t^{\Re(w)}\Big[1+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big],\quad\forall t\geq 1.$ (3)
Let us introduce some notations.
Notation 1.
For every $\eta\in L^{\infty}(\mathbb{C})$ and $w\in\mathbb{C}$, we denote by $\psi_{\eta,w}(\cdot)$ the unique continuous solution of the differential equation (2) given by equation (3).
Notation 2.
Denote by $\mathbb{C}_{+}$ the right half-plane defined as
$\mathbb{C}_{+}:=\Big\{w\in\mathbb{C}:\quad\Re(w)>0\Big\}.$
For every $\eta\in L^{\infty}(\mathbb{C})$, let $\mu_{\eta}$ denote the function $\mu_{\eta}:\mathbb{C}_{+}\to\mathbb{C}$, defined as
$\mu_{\eta}(w)=-1-(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall w\in\mathbb{C}_{+}.$
The function $\mu_{\eta}$ is defined for all $s\in\mathbb{C}_{+}$. Indeed, the integral is absolutely convergent, since $\eta\in L^{\infty}(\mathbb{C})$ and since for all $w\in\mathbb{C}_{+}$ we have $\Re(w)>0$. In the following lemma, we are interested in bounded solutions on $1,+\infty)$ and their initial condition.
Lemma 3.
Let $\eta\in L^{\infty}(\mathbb{C})$ and $w\in\mathbb{C}_{+}$. Consider the continuous solution $\psi_{\eta,w}(\cdot)$ of the differential equation (2), as introduced in Notation 1. Then,
$\sup_{t\geq 1}\left|\psi_{\eta,w}(t)\right|<+\infty\quad\text{if and only if}\quad\mu_{\eta}(w)=0,$
where the function $\mu_{\eta}$ is defined in Notation 2. Furthermore, if $\eta$ satisfies the rotation number hypothesis (H), then
$\sup_{t\geq 1}\left|t\left(\psi_{\eta,w}\big(t\big)-\rho\frac{w-1}{w}t^{-i\Im(w)}\right)\right|<+\infty,$
where $\rho$ denotes the rotation number of $\eta$.
Proof.
Let $w\in\mathbb{C}_{+}$ be fixed. Suppose that $\sup_{t\geq 1}|\psi_{\eta,w}(t)|<+\infty$. By equation (3), this implies
$\sup_{t\geq 1}\Bigg[t^{\Re(w)}\Big|1+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big|\Bigg]<+\infty.$
Since $\Re(w)>0$, we must have
$\Big|1+(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du\Big|=0,$
which gives $\mu_{\eta}(w)=0$. Now, prove that $\sup_{t\geq 1}|\psi_{\eta,w}(t)|<+\infty$. By definition of $\mu_{\eta}(w)$, we have
$\mu_{\eta}(w)+1+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du=-(1-w)\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\ \forall t\geq 1.$
Suppose that $\mu_{\eta}(w)=0$, using equation (3), we obtain
$\psi_{\eta,w}(t)=-(1-w)t^{\Re(w)}\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$ (4)
By Hypothesis $\eta\in L^{\infty}(\mathbb{C})$ then there exists a constant $c>0$ such that $|\eta(t)|<c$ almost everywhere, we obtain
$\sup_{t\geq 1}|\psi_{\eta,w}(t)|\leq c\frac{|1-w|}{\Re(w)}<+\infty.$
Now, prove the second item of the present Lemma. Suppose that the function $\eta$ satisfies the hypothesis (H). Then there exists $\tilde{c}>0$ such that
$\Big|\int_{1}^{t}\,\big(\eta(u)-\rho\big)du\Big|<\tilde{c},\quad\forall t\geq 1,$
where $\rho$ is the rotation number of $\eta$. Hence, the equation (4) can be written as
$\psi_{\eta,w}(t)=\rho\frac{w-1}{w}t^{-i\,\Im(w)}-(1-w)t^{{\Re(w)}}\int_{t}^{+\infty}u^{-1-w}(\eta(u)-\rho)du,$
Using the integration by parts formula, we obtain
$\psi_{\eta,w}(t)=\rho\frac{w-1}{w}t^{-i\,\Im(w)}-(1-w^{2})t^{\Re(w)}\int_{t}^{+\infty}u^{-2-w}\int_{t}^{u}(\eta(v)-\rho)dvdu.$
For every $t\geq 1$, we get
$\left|\psi_{\eta,w}(t)-\rho\frac{w-1}{w}t^{-i\,\Im(w)}\right|<2\tilde{c}\frac{|1-w^{2}|}{1+\Re(w)}t^{-1}.$
We denote by $B\subset\mathbb{C}$ the critical strip, excluding the critical line and the real axis, defined as
$B:=\Big\{w\in\mathbb{C}:\Re(w)\in(0,1),\ \Re(w)\neq\tfrac{1}{2},\ \Im(w)\neq 0\Big\}.$
The following main result provides an answer to the Dynamical Conjecture formulated in [1]. We shows that for a function $\eta$ satisfying the Rotation number hypothesis (H), for every $s\in B$, the continuous solutions $\psi_{\eta,s}(\cdot)$ and $\psi_{\eta,1-\overline{s}}(\cdot)$ of the differential equation (2) cannot both be bounded on $1,+\infty)$.
Theorem 4.
Let $\eta\in L^{\infty}(\mathbb{C})$ satisfy the Rotation Number Hypothesis (H). Suppose that the function $\mu_{\eta}$ satisfies:
$\forall\beta\in\mathbb{R}^{*}:\quad\mu_{\eta}(1+i\,\beta)\neq 0.$
Then, for every $s\in B$, we have
$\big(\mu_{\eta}(s),\,\mu_{\eta}(1-\overline{s})\big)\neq(0,0).$
2 Main proposition
Let us introduce the following notation.
Notation 5.
We denote by $I\subset\mathbb{C}$ the following set
$I:=\Big\{w\in\mathbb{C}:\Re(w)\in[0,1],\ \Re(w)\neq\tfrac{1}{2},\ \Im(w)\neq 0\Big\}.$
Consider the continuous solutions of the differential equation (2) as given in Notation 1. For every $w\in I$, we denote by $\delta_{w}$ the function
$\delta_{\eta,w}(t):=\frac{1}{2\Re(w)-1}\Big(\psi_{\eta,w}(t)-\psi_{\eta,1-\overline{w}}(t)\Big),\quad\forall t\geq 1.$
The differential equation (2) can be viewed as a differential equation with the parameter $\Re(w)$. Intuitively, this introduces an order structure within the set of solutions with a same initial condition and parameterized by $\Re(w)$. We start to exhibit this order between solutions through the following proposition.
Proposition 6.
For every $s:=\sigma+i\tau\in I$, the function $\delta_{\eta,s}$, defined in Notation (5), satisfies the following integral equation:
$t^{-\frac{1}{2}}\delta_{\eta,s}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
$+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
$+\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\ du.$
Proof.
Let $w\in\mathbb{C}_{+}$ and denote $w:=\alpha+i\,\beta$. For every $t\geq 1$ denote
$x_{w}(t):=t^{-\frac{1}{2}}\psi_{\eta,w}(t).$ (5)
According to Notation 1, we obtain the following differential equation
$\dot{x}_{w}(t)=\big(\alpha-\frac{1}{2}\big)\,t^{-1}x_{w}(t)+(1-w)\,t^{-\frac{3}{2}-i\,\beta}\,\eta(t),\quad x_{w}(1)=1.$ (6)
Let $s:=\sigma+i\,\tau\in I$. From the previous equation, we have
$\dot{x_{s}}(t)+\dot{x}_{1-\overline{s}}(t)=(\sigma-\frac{1}{2})t^{-1}\big({x_{s}}(t)-{x}_{1-\overline{s}}(t)\big)+(1-2i\tau)t^{-\frac{3}{2}-i\,\tau}\,\eta(t).$
Use the fact that $x_{s}(1)=x_{1-\overline{s}}(1)=1$ and integrate, we obtain
${x_{s}}(t)+{x}_{1-\overline{s}}(t)$ $=(\sigma-\frac{1}{2})\int_{1}^{t}u^{-1}\big({x_{s}}(u)-{x}_{1-\overline{s}}(u)\big)\,du$
$+2+(1-2i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\,du.$
From the differential equation (6), we have
$\dot{x}_{s}(t)-\dot{x}_{1-\overline{s}}(t)$ $=(\sigma-\frac{1}{2})t^{-1}\big(x_{s}(t)+x_{1-\overline{s}}(t)\big)$
$-(2\sigma-1)t^{-\frac{3}{2}-i\,\tau}\,\eta(t).$
The two previous differential equations, gives
$\dot{x}_{s}(t)-\dot{x}_{1-\overline{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{-1}\int_{1}^{t}u^{-1}\big({x_{s}}(u)-{x}_{1-\overline{s}}(u)\big)\,du$
$+(2\sigma-1)t^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\,du\Big]$
$+(2\sigma-1)t^{-\frac{3}{2}-i\,\tau}\,\eta(t).$
Use the fact that ${x}_{s}(1)-{x}_{1-\overline{s}}(1)=0$ and integrate, we obtain
${x}_{s}(t)-{x}_{1-\overline{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-1}\big({x_{s}}(v)-{x}_{1-\overline{s}}(v)\big)\,dv\,du$
$+(2\sigma-1)\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
$+(2\sigma-1)\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u),\ du.$ (7)
By the notation (5), we have
${x_{s}}(t)-{x}_{1-\overline{s}}(t)=t^{-\frac{1}{2}}\big({\psi_{\eta,s}}(t)-{\psi}_{\eta,1-\overline{s}}(t)\big).$
By the Notation (5), we have
$\delta_{\eta,s}(t)=\frac{1}{2\sigma-1}\big({\psi_{\eta,s}}(t)-{\psi}_{\eta,1-\overline{s}}(t)\big).$
Then
${x_{s}}(t)-{x}_{1-\overline{s}}(t)=(2\sigma-1)t^{-\frac{1}{2}}\delta_{\eta,s}(t),$
In term of $\delta_{s}$ the integral equation (7) becomes
$t^{-\frac{1}{2}}\delta_{\eta,s}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
$+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
$+\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\ du.$
3 Proof of the Theorem 4
Lemma 7.
For every $s\in I$ with $\Re(s)=\sigma$, the function $\delta_{\eta,s}$, defined in Notation (5), satisfies the following Volterra integral equation:
$\delta_{\eta,s}(t)=\delta_{\eta,1+i\Im(s)}(t)-\frac{\sigma(1-\sigma)}{2\sigma-1}\int_{1}^{t}\left[\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}\right]v^{-1}\delta_{\eta,1+i\Im(s)}(v)\,dv.$
Proof.
By the Proposition 6 every $s\in I$, we have
$t^{-\frac{1}{2}}\delta_{\eta,s}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
$+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
$+\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\ du.$
By definition of the set II in Notation 5, we have in particular,
$t^{-\frac{1}{2}}\delta_{\eta,1+i\Im(s)}(t)=$ $\frac{1}{4}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,1+i\Im(s)}(v)\,dv\,du$
$+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
$+\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\ du.$
Subtracting the two previous equations gives:
$\delta_{\eta,s}(t)-\delta_{\eta,1+i\Im(s)}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
$-\frac{1}{4}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,1+i\Im(s)}(v)\,dv\,du.$
After two differentiations, to the second‑order linear differential equation
$\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\Im(s)}(t)\Big)=\sigma(\sigma-1)t^{-2}\delta_{\eta,s}(t),$
with initial conditions $\delta_{\eta,s}(1)=\delta_{\eta,1+i\Im(s)}(1)$ and $\delta_{\eta,s}^{\prime}(1)=\delta_{\eta,1+i\Im(s)}^{\prime}(1)$. Solving this Cauchy problem gives the explicit representation
$\delta_{\eta,s}(t)=\delta_{\eta,1+i\Im(s)}(t)-\frac{\sigma(1-\sigma)}{2\sigma-1}\int_{1}^{t}\left[\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}\right]v^{-1}\delta_{\eta,1+i\Im(s)}(v)\,dv.$
Proof of the Theorem 4.
Let $s\in B$ and $\sigma=\Re(s)$. By the Lemma 7, we have
$\delta_{\eta,s}(t)$ $=\delta_{\eta,1+i\Im(s)}(t)$ (8)
$-\frac{\sigma(1-\sigma)}{2\sigma-1}\int_{1}^{t}\left[\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}\right]v^{-1}\delta_{\eta,1+i\Im(s)}(v)\,dv.$
By equation (3), we have
$\delta_{\eta,1+i,\Im(s)}(t)$ $=t\Big[1-i\ \Im(s)\int_{1}^{t}u^{-2-i\Im(s)}\eta(u)\,du\Big]$
$-1-(1-i\ \Im(s))\int_{1}^{t}u^{-1-i\Im(s)}\eta(u)\,du,\quad\forall t\geq 1,$
which can be written as
$\delta_{\eta,1+i\Im(s)}(t)$ $=-t\mu\big(1+i\Im(s)\big)+i\ \Im(s)t\int_{t}^{+\infty}u^{-2-i\Im(s)}\eta(u)\,du$
$-1-\rho\frac{1-i\ \Im(s)}{i\Im(s)}\big(1-t^{-i\Im(s)}\big)$
$-(1-i\ \Im(s))\int_{1}^{t}u^{-1-i\Im(s)}\big(\eta(u)-\rho\big)\,du,\quad\forall t\geq 1.$
Applying the integration by parts formula to the last integral of the previous equation and using the hypothesis (H), there exists $M_{\Im(s)}>0$ such that
$\Big|\delta_{\eta,1+i\Im(s)}(t)+t\mu\big(1+i\Im(s)\big)\Big|<M_{\Im(s)},\quad\forall t>1.$ (9)
By hypothesis, $\mu_{\eta}(1+i\beta)\neq 0$. Assume $\Im(\mu_{\eta}(1+i\beta))\neq 0$ (the real case is analogous). Equation (9) implies that for sufficiently large $t$, $\Im(\delta_{\eta,1+i\Im(s)}(t))$ maintains a constant sign and grows linearly. It follows from (8) that:
$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}\to 0^{+}\text{ as }t\to+\infty.$ (10)
In fact
$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}=1-\frac{\sigma(1-\sigma)}{2\sigma-1}I(t),$
where
$I(t)=\int_{1}^{t}\left[\frac{\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}\right]v^{-1}\,\Im\big(\delta_{\eta,1+i\Im(s)}(v)\big)\,dv$
From estimate (9), as $t\to+\infty$ we have
$\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)=-\,t\Im\big(\mu_{\eta}(1+i\tau)\big)+O(1).$
Hence
$\frac{\Im\big(\delta_{\eta,1+i\tau}(v)\big)}{\Im\big(\delta_{\eta,1+i\tau}(t)\big)}=\frac{v}{t}+o(1)$
uniformly for $1\leq v\leq t$. Substituting this relation into the integral kernel gives
$I(t)\sim\int_{1}^{t}\left[t^{\sigma}v^{-\sigma}-t^{1-\sigma}v^{\sigma-1}\right]v^{-1}\frac{v}{t}\,dv.$
Thus
$I(t)\sim\frac{1}{t}\int_{1}^{t}\left(t^{\sigma}v^{-\sigma}-t^{1-\sigma}v^{\sigma-1}\right)dv.$
We compute
$\int_{1}^{t}t^{\sigma}v^{-\sigma}dv=t^{\sigma}\frac{t^{1-\sigma}-1}{1-\sigma}=\frac{t-t^{\sigma}}{1-\sigma},$
and
$\int_{1}^{t}t^{1-\sigma}v^{\sigma-1}dv=t^{1-\sigma}\frac{t^{\sigma}-1}{\sigma}=\frac{t-t^{1-\sigma}}{\sigma}.$
Therefore
$I(t)\sim\frac{1}{t}\left[\frac{t-t^{\sigma}}{1-\sigma}-\frac{t-t^{1-\sigma}}{\sigma}\right].$
After simplification we obtain
$I(t)\sim\frac{(2\sigma-1)t-\sigma t^{\sigma}+(1-\sigma)t^{1-\sigma}}{\sigma(1-\sigma)t}.$
Substituting into the previous expression yields
$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\tau}(t)\big)}\sim 1-\frac{\sigma(1-\sigma)}{2\sigma-1}\frac{(2\sigma-1)t-\sigma t^{\sigma}+(1-\sigma)t^{1-\sigma}}{\sigma(1-\sigma)t}.$
After simplification we obtain
$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\tau}(t)\big)}\sim\frac{\sigma}{(2\sigma-1)t^{1-\sigma}}-\frac{1-\sigma}{(2\sigma-1)t^{\sigma}}>0\quad\text{as}\quad t\to+\infty.$
If $\mu_{\eta}(s)=\mu_{\eta}(1-\overline{s})=0$, Lemma 3 implies that there exists $C_{s}>0$ such that
$\left|\Im\big(\delta_{\eta,s}(t)\big)-\Im\left(\rho\left(\frac{s-1}{s}+\frac{\overline{s}}{1-\overline{s}}\right)t^{-i\Im(s)}\right)\right|<C_{s}t^{-1},\quad\forall t\geq 1.$
Since $s\in B$, the $\Re(s)\neq\frac{1}{2}$ and $\Im(s)\neq 0$. Use the fact that $\rho\neq 0$, then
$\rho\big(\frac{s-1}{s}+\frac{\overline{s}}{1-\overline{s}}\big)\neq 0.$
Hence, there exists a sequence $(t_{n})_{n\in\mathbb{N}}\subset[1,+\infty)$ such that $\lim_{n\to+\infty}t_{n}=+\infty$ and such that
$\Im\big(\delta_{\eta,s}(t_{n})\big)=0,\quad\forall n\in\mathbb{N},$
From equation (9) and the assumption $\Im(\mu_{\eta}(1+i\beta))\neq 0$, there exists $n_{s}\in\mathbb{N}$ such that
$\frac{\Im\big(\delta_{\eta,s}(t_{n})\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t_{n})\big)}=0,\quad\forall n\geq n_{s},$
In other words, the function
$t\mapsto\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}$
oscillates like $t^{-1}\Im\big(\omega_{s}t^{-i\Im(s)}\big)$ with $\omega_{s}\in\mathbb{C}^{*}$, preventing the ratio in (10) from remaining strictly positive or converging as required, thus yielding a contradiction. ∎