### 15 Apr 2026 v3

# Rigidity and Structural Asymmetry of Bounded Solutions

Abstract
In this manuscript, we introduce a family of parametrized non-homogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter. We identify a Rotation number hypothesis on the non-homogeneous term, which establishes a structural asymmetry: if two solutions with the same initial condition equal to 11, corresponding respectively to the parameters $s$ and $1-\overline{s}$ lying in the critical strip, are both bounded on $1,+\infty)$, then $\Re(s)=\tfrac{1}{2}$.
Keywords: Non-homogeneous complex linear differential equation, bounded solutions, Euler differential equation, Volterra integral equation.
MSC subject classifications: 34A30, 34E05
1 Introduction
In this manuscript, we study the following non-homogeneous linear complex differential equations:
$\dot{\theta}=wt^{-1}\theta+t^{-1}\eta(t),\quad\theta(1)=\frac{1}{1-w},\quad\theta:[1,+\infty)\to\mathbb{C},$ (1)
where $w\in\mathbb{C}$ such that $\Im(w)\neq 0$ is the parameter and the non-homogeneous term $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ where $\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ is the set of bounded and locally integrable functions from $1,+\infty)$ to $\mathbb{C}$. We are interested in the initial conditions of the bounded solutions of the previous differential equation. To this end, we analyze the behavior of the transformed solution $t\mapsto(1-w)t^{-i\Im(w)}\theta(t)$, which leads us to consider the following differential equation:
$\dot{x}=\Re(w)t^{-1}x+(1-w)t^{-1-i\Im(w)}\eta(t),$ (2)
$t\in[1,+\infty),\quad x(1)=1,\quad x:[1,+\infty)\to\mathbb{C}.$
Since $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$, the function
$t\mapsto t^{\Re(w)}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1.$
is absolutely continuous on $1,+\infty)$. The differential equation (2) is a non-homogeneous linear differential equation. Then, there exists a unique continuous solution $\psi_{\eta,w}(\cdot)$ of (2) such that $\psi_{w}(1)=1$, which is given by
$\psi_{\eta,w}(t)=t^{\Re(w)}\Big[1+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big],\quad\forall t\geq 1.$ (3)
Let us introduce some notations.
Notation 1.
For every $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ and $w\in\mathbb{C}$, we denote by $\psi_{\eta,w}(\cdot)$ the unique continuous solution of the differential equation (2) given by equation (3).
Notation 2.
Denote by $\mathbb{C}_{+}$ the right half-plane defined as
$\mathbb{C}_{+}:=\Big\{w\in\mathbb{C}:\quad\Re(w)>0\Big\}.$
For every $f\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$, let $\mu_{f}$ denote the function $\mu_{f}:\mathbb{C}_{+}\to\mathbb{C}$, defined as
$\mu_{f}(w)=-1-(1-w)\int_{1}^{+\infty}u^{-1-w}f(u)\,du,\quad\forall w\in\mathbb{C}_{+}.$
The function $\mu_{f}$ is defined for all $s\in\mathbb{C}_{+}$. Indeed, the integral is absolutely convergent, since $f\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ and since for all $w\in\mathbb{C}_{+}$ we have $\Re(w)>0$. In the following lemma, we are interested in bounded solutions on $1,+\infty)$ and their initial condition.
Lemma 3.
Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ and $w\in\mathbb{C}_{+}$. Consider the continuous solution $\psi_{\eta,w}(\cdot)$ of the differential equation (2), as introduced in Notation 1. Then,
$\sup_{t\geq 1}\left|\psi_{\eta,w}(t)\right|<+\infty\quad\text{if and only if}\quad\mu_{\eta}(w)=0,$
where the function $\mu_{\eta}$ is defined in Notation 2.
Proof.
Let $w\in\mathbb{C}_{+}$ be fixed. Suppose that $\sup_{t\geq 1}|\psi_{\eta,w}(t)|<+\infty$. By equation (3), this implies
$\sup_{t\geq 1}\Bigg[t^{\Re(w)}\Big|1+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big|\Bigg]<+\infty.$
Since $\Re(w)>0$, we must have
$\Big|1+(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du\Big|=0,$
which gives $\mu_{\eta}(w)=0$. Now, prove that $\sup_{t\geq 1}|\psi_{\eta,w}(t)|<+\infty$. By definition of $\mu_{\eta}(w)$, we have
$\mu_{\eta}(w)+1+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du=-(1-w)\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\ \forall t\geq 1.$
Suppose that $\mu_{\eta}(w)=0$, using equation (3), we obtain
$\psi_{\eta,w}(t)=-(1-w)t^{\Re(w)}\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$ (4)
By Hypothesis $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ then there exists a constant $c>0$ such that $|\eta(t)|<c$ for all $t\geq 1$, we obtain
$\sup_{t\geq 1}|\psi_{\eta,w}(t)|\leq c\frac{|1-w|}{\Re(w)}<+\infty.$
2 Main result
As formulated in the Dynamical Conjecture in [1], we assume that the function $\eta$ satisfies the following Rotation number hypothesis:
$\exists\rho_{\eta}\in\mathbb{C}^{*}:\quad\sup_{t\geq 1}\Big|\int_{1}^{t}\,\big(\eta(u)-\rho_{\eta}\big)du\Big|<+\infty,$ (H)
The complex number $\rho_{\eta}$ is called the rotation number of $\eta$. Notice that if $\rho_{\eta}$ exists, then it is unique; this can be proved by contradiction. Denote by $\mathcal{L}[\eta]$ the following transformation
$\mathcal{L}[\eta](t):=\int_{1}^{t}\,\big(\eta(u)-\rho_{\eta}\big)du,\quad\forall t\geq 1.$
The Rotation number hypothesis implies that $\mathcal{L}[\eta]\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$. Hence, the function $\mu_{\mathcal{L}[\eta]}$ is defined on $\mathbb{C}_{+}$ as given in the Notation 2.
We denote by $B\subset\mathbb{C}$ the critical strip, excluding the critical line and the real axis, defined as
$B:=\Big\{w\in\mathbb{C}:\Re(w)\in(0,1),\ \Re(w)\neq\tfrac{1}{2},\ \Im(w)\neq 0\Big\}.$
The following main result provides an answer to the Dynamical Conjecture formulated in [1]. We shows that for a function $\eta$ satisfying the Rotation number hypothesis (H), for every $s\in B$, the continuous solutions $\psi_{\eta,s}(\cdot)$ and $\psi_{\eta,1-\overline{s}}(\cdot)$ of the differential equation (2) cannot both be bounded on $1,+\infty)$.
Theorem 4.
Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ satisfy the Rotation Number Hypothesis (H). Suppose that the functions $\mu_{\eta}$ and $\mathcal{L}[\eta]$ satisfies:
$\forall x\in\mathbb{R}^{*}:\quad\mu_{\eta}(1+i\,x)\neq 0\quad\text{and}\quad\mu_{\mathcal{L}[\eta]}(1+i\,x)\neq\frac{\rho_{\eta}-ix}{1+ix}-\frac{1}{|1+i\,x|^{2}}.$
Then, for every $s\in B$, we have
$\big(\mu_{\eta}(s),\,\mu_{\eta}(1-\overline{s})\big)\neq(0,0).$
Remark 5.
In the proof of the Lemma 9, we note that
$\frac{|1+i\,x|^{2}}{ix}\Big[\frac{\rho_{\eta}-ix}{1+ix}-\mu_{\mathcal{L}[\eta]}(1+i\,x)-\frac{1}{|1+i\,x|^{2}}\Big]=\lim_{r\to_{>}0}\mu_{\eta}(r+i\,x),$
which is the analytic extension of the function $\mu_{\eta}$ to the line $\Re(z)=0$. Consequently, the condition of the previous theorem is equivalent to
$\forall x\in\mathbb{R}^{*}:\quad\mu_{\eta}(1+i\,x)\neq 0\quad\text{and}\quad\mu_{\eta}(i\,x)\neq 0.$
3 Main proposition
Let us introduce the following notation.
Notation 6.
We denote by $I\subset\mathbb{C}$ the following set
$I:=\Big\{w\in\mathbb{C}:\Re(w)\in[0,1],\ \Re(w)\neq\tfrac{1}{2},\ \Im(w)\neq 0\Big\}.$
Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$. Consider the continuous solutions of the differential equation (2) as given in Notation 1. For every $w\in I$, we denote by $\delta_{w}$ the function
$\delta_{\eta,w}(t):=\frac{1}{2\Re(w)-1}\Big(\psi_{\eta,w}(t)-\psi_{\eta,1-\overline{w}}(t)\Big),\quad\forall t\geq 1.$
The differential equation (2) can be viewed as a differential equation with the parameter $\Re(w)$. Intuitively, this introduces an order structure within the set of solutions with a same initial condition and parameterized by $\Re(w)$. We start to exhibit this order between solutions through the following proposition.
Proposition 7.
Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$. For every $s:=\sigma+i\tau\in I$, the functions $\delta_{\eta,s}$ and $\delta_{\eta,1+i\tau}$, defined in Notation (6), satisfies the following Volterra integral equation:
$\delta_{\eta,1+i\tau}(t)$ $=\sigma(1-\sigma)t^{1-\sigma}\int_{1}^{t}\,u^{-2(1-\sigma)}\int_{1}^{u}\,v^{-1-\sigma}\delta_{\eta,1+i\tau}(v)\,dv\ du$
$+\delta_{\eta,s}(t),\quad\forall t\geq 1,$
Lemma 8.
Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$. For every $s:=\sigma+i\tau\in I$, the functions $\delta_{\eta,s}$ and $\delta_{\eta,1+i\tau}$, defined in Notation (6), satisfies the following Euler differential equation:
$\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\Big)=-\sigma(1-\sigma)t^{-2}\delta_{\eta,s}(t),\quad\forall t\geq 1,$
with initial conditions $\delta_{\eta,s}(1)=\delta_{\eta,1+i\tau}(1)=0$ and $\dot{\delta}_{\eta,s}(1)-\dot{\delta}_{\eta,1+i\tau}(1)=0$.
Proof.
Let $w\in\mathbb{C}_{+}$ and denote $w:=\alpha+i\,\beta$. For every $t\geq 1$ denote
$x_{w}(t):=t^{-\frac{1}{2}}\psi_{\eta,w}(t).$ (5)
According to Notation 1, we obtain the following differential equation
$\dot{x}_{w}(t)=\big(\alpha-\frac{1}{2}\big)\,t^{-1}x_{w}(t)+(1-w)\,t^{-\frac{3}{2}-i\,\beta}\,\eta(t),\quad x_{w}(1)=1.$ (6)
Let $s:=\sigma+i\,\tau\in I$. From the previous equation, we have
$\dot{x_{s}}(t)+\dot{x}_{1-\overline{s}}(t)=(\sigma-\frac{1}{2})t^{-1}\big({x_{s}}(t)-{x}_{1-\overline{s}}(t)\big)+(1-2i\tau)t^{-\frac{3}{2}-i\,\tau}\,\eta(t).$
Use the fact that $x_{s}(1)=x_{1-\overline{s}}(1)=1$ and integrate, we obtain
${x_{s}}(t)+{x}_{1-\overline{s}}(t)$ $=(\sigma-\frac{1}{2})\int_{1}^{t}u^{-1}\big({x_{s}}(u)-{x}_{1-\overline{s}}(u)\big)\,du$
$+2+(1-2i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\,du.$
From the differential equation (6), we have
$\dot{x}_{s}(t)-\dot{x}_{1-\overline{s}}(t)$ $=(\sigma-\frac{1}{2})t^{-1}\big(x_{s}(t)+x_{1-\overline{s}}(t)\big)$
$-(2\sigma-1)t^{-\frac{3}{2}-i\,\tau}\,\eta(t).$
The two previous differential equations, gives
$\dot{x}_{s}(t)-\dot{x}_{1-\overline{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{-1}\int_{1}^{t}u^{-1}\big({x_{s}}(u)-{x}_{1-\overline{s}}(u)\big)\,du$
$+(2\sigma-1)t^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\,du\Big]$
$+(2\sigma-1)t^{-\frac{3}{2}-i\,\tau}\,\eta(t).$
Use the fact that ${x}_{s}(1)-{x}_{1-\overline{s}}(1)=0$ and integrate, we obtain
${x}_{s}(t)-{x}_{1-\overline{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-1}\big({x_{s}}(v)-{x}_{1-\overline{s}}(v)\big)\,dv\,du$
$+(2\sigma-1)\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
$+(2\sigma-1)\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u),\ du.$ (7)
By the notation (5), we have
${x_{s}}(t)-{x}_{1-\overline{s}}(t)=t^{-\frac{1}{2}}\big({\psi_{\eta,s}}(t)-{\psi}_{\eta,1-\overline{s}}(t)\big).$
By the Notation (6), we have
$\delta_{\eta,s}(t)=\frac{1}{2\sigma-1}\big({\psi_{\eta,s}}(t)-{\psi}_{\eta,1-\overline{s}}(t)\big).$
Then
${x_{s}}(t)-{x}_{1-\overline{s}}(t)=(2\sigma-1)t^{-\frac{1}{2}}\delta_{\eta,s}(t),$
In term of $\delta_{s}$ the integral equation (7) becomes
$t^{-\frac{1}{2}}\delta_{\eta,s}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
$+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
$+\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\ du.$
By definition of the set II in Notation 6, we have in particular,
$t^{-\frac{1}{2}}\delta_{\eta,1+i\tau}(t)=$ $\frac{1}{4}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,1+i\tau}(v)\,dv\,du$
$+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\tau)\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\Big]du$
$+\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\ du.$
Subtracting the two previous equations gives:
$\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
$-\frac{1}{4}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,1+i\tau}(v)\,dv\,du.$
After two differentiations, to the second‑order linear differential equation
$\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\Big)=-\sigma(1-\sigma)t^{-2}\delta_{\eta,s}(t),$
with initial conditions $\delta_{\eta,s}(1)=\delta_{\eta,1+i\tau}(1)=0$ and $\dot{\delta}_{\eta,s}(1)-\dot{\delta}_{\eta,1+i\tau}(1)=0$. ∎
Proof of the Proposition 7.
By the Lemma 8 every $s:=\sigma+i\,\tau\in I$, we have
$\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\Big)=-\sigma(1-\sigma)t^{-2}\delta_{\eta,s}(t),$
which can be written as
$t^{1-\sigma}\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\Big)=-\sigma(1-\sigma)t^{-1-\sigma}\delta_{\eta,s}(t),$
Use the fact that $\delta_{\eta,s}(1)-\delta_{\eta,1+i\tau}(1)=0$ and $\dot{\delta}_{\eta,s}(1)-\dot{\delta}_{\eta,1+i\tau}(1)=0$ and use the integration by part formula,
$t^{1-\sigma}\frac{d}{dt}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\big)$ $-(1-\sigma)\int_{1}^{t}\,u^{-\sigma}\frac{d}{du}\big(\delta_{\eta,s}(u)-\delta_{\eta,1+i\tau}(u)\big)\,du$
$=-\sigma(1-\sigma)\int_{1}^{t}\,u^{-1-\sigma}\delta_{\eta,s}(u)\,du,$
$t^{1-\sigma}\frac{d}{dt}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\big)$ $-(1-\sigma)t^{-\sigma}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\big)$
$-\sigma(1-\sigma)\int_{1}^{t}\,u^{-1-\sigma}\big(\delta_{\eta,s}(u)-\delta_{\eta,1+i\tau}(u)\big)\,du$
$=-\sigma(1-\sigma)\int_{1}^{t}\,u^{-1-\sigma}\delta_{\eta,s}(u)\,du,$
We obtained
$t^{1-\sigma}\frac{d}{dt}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\big)$ $-(1-\sigma)t^{-\sigma}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\big)$
$+\sigma(1-\sigma)\int_{1}^{t}\,u^{-1-\sigma}\delta_{\eta,1+i\tau}(u)du$
$=0,\quad\forall t\geq 1.$
That can be written as
$\frac{d}{dt}\Big(t^{-1+\sigma}\big(\delta_{\eta,1+i\tau}(t)-\delta_{\eta,s}(t)\big)\Big)$ $=\sigma(1-\sigma)t^{-2(1-\sigma)}\int_{1}^{t}\,u^{-1-\sigma}\delta_{\eta,1+i\tau}(u)du$
$+\frac{d}{dt}\big(t^{-1+\sigma}\delta_{\eta,s}(t)\big),\quad\forall t\geq 1,$
Integrate
$\delta_{\eta,1+i\tau}(t)$ $=\sigma(1-\sigma)t^{1-\sigma}\int_{1}^{t}\,u^{-2(1-\sigma)}\int_{1}^{u}\,v^{-1-\sigma}\delta_{\eta,1+i\tau}(v)\,dv\ du$
$+\delta_{\eta,s}(t),\quad\forall t\geq 1,$
4 Proof of the Theorem 4
Lemma 9.
Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ satisfy the Rotation Number Hypothesis (H). For every $x\in\mathbb{R}^{*}$, the function $\delta_{\eta,1+ix}$, defined in Notation (6), satisfies the following equation
$\sup_{t\geq 1}\Big|\int_{1}^{t}\,u^{-1}\delta_{\eta,1+ix}(u)\,du+\alpha_{x}t-\beta_{x}\ln(t)\Big|<+\infty,$
where
$\alpha_{x}:=\mu_{\eta}\big(1+ix\big)\ \text{and}\ \beta_{x}:=\frac{|1+i\,x|^{2}}{ix}\Big[\frac{\rho_{\eta}-ix}{1+ix}-\mu_{\mathcal{L}[\eta]}(1+i\,x)-\frac{1}{|1+i\,x|^{2}}\Big].$
Proof.
Let $x\in\mathbb{R}^{*}$. By the Notation 6 and equation (3), we have
$\delta_{\eta,1+i,x}(t)$ $=t\Big[1-i\,x\int_{1}^{t}u^{-2-ix}\eta(u)\,du\Big]$
$-1-(1-i\,x)\int_{1}^{t}u^{-1-ix}\eta(u)\,du,\quad\forall t\geq 1,$
which can be written as
$\delta_{\eta,1+ix}(t)$ $=-t\mu_{\eta}\big(1+ix\big)+i\,xt\int_{t}^{+\infty}u^{-2-ix}\eta(u)\,du$
$-1-\rho\frac{1-i\,x}{ix}\big(1-t^{-ix}\big)$
$-(1-i\,x)\int_{1}^{t}u^{-1-ix}\big(\eta(u)-\rho\big)\,du,\quad\forall t\geq 1.$
Use the Rotation number hypothesis (H) and use the integration by formula
$\delta_{\eta,1+ix}(t)$ $=-t\mu_{\eta}\big(1+ix\big)-\rho\frac{1-i\,x}{ix}\big(1-t^{-ix}\big)+\rho\frac{i\,x}{1+ix}t^{-ix}$
$-1-|1+i\,x|^{2}\int_{1}^{+\infty}u^{-2-ix}\int_{1}^{u}\,\big(\eta(v)-\rho\big)\,dv\,du$
$+i\,xt\int_{t}^{+\infty}u^{-2-ix}\big(\eta(u)-\rho\big)\,du$
$+(1-i\,x)\int_{t}^{+\infty}u^{-1-ix}\big(\eta(u)-\rho\big)\,du,\quad\forall t\geq 1.$
Equivalent to,
$\delta_{\eta,1+ix}(t)$ $=-t\mu_{\eta}\big(1+ix\big)-\rho\frac{1-i\,x}{ix}\big(1-t^{-ix}\big)+\rho\frac{i\,x}{1+ix}t^{-ix}$
$-1-|1+i\,x|^{2}\frac{1}{ix}\big(1+\mu_{\mathcal{L}[\eta]}(1+i\,x)\big)$
$+i\,xt\int_{t}^{+\infty}u^{-2-ix}\big(\eta(u)-\rho\big)\,du$
$+(1-i\,x)\int_{t}^{+\infty}u^{-1-ix}\big(\eta(u)-\rho\big)\,du,\quad\forall t\geq 1.$
or,
$\delta_{\eta,1+ix}(t)$ $=-t\alpha_{x}+\beta_{x}$
$-\rho\frac{1}{ix(1+ix)}t^{-ix}$
$+i\,x(2+ix)t\int_{t}^{+\infty}u^{-3-ix}\int_{t}^{u}\,\big(\eta(v)-\rho\big)\,dv\,\,du$
$+|1+i\,x|^{2}\int_{t}^{+\infty}u^{-2-ix}\int_{t}^{u}\,\big(\eta(v)-\rho\big)\,dv\,du,\quad\forall t\geq 1.$
where
$\alpha_{x}:=\mu_{\eta}\big(1+ix\big)\ \text{and}\ \beta_{x}:=\frac{|1+i\,x|^{2}}{ix}\Big[\frac{\rho_{\eta}-ix}{1+ix}-\mu_{\mathcal{L}[\eta]}(1+i\,x)-\frac{1}{|1+i\,x|^{2}}\Big].$
Using the hypothesis (H), we have
$\sup_{u\geq t\geq 1}\Big|\int_{t}^{u}\,\big(\eta(v)-\rho\big)\,dv\Big|<+\infty,$
Then
$\sup_{t\geq 1}\Big|\int_{1}^{t}\,u^{-1}\delta_{\eta,1+ix}(u)\,du+\alpha_{x}t-\beta_{x}\ln(t)\Big|<+\infty,$
Proof of the Theorem 4.
Let $s:=\sigma+i\,\tau\in B$. By the Lemma 9, there exists $M_{\tau}>0$ such that
$\Big|\int_{1}^{t}\,u^{-1}\delta_{\eta,1+i\tau}(u)\,du+\alpha_{\tau}t-\beta_{\tau}\ln(t)\Big|<M_{\tau},\quad\forall t\geq 1,$ (8)
where by the hypothesis of the present Theorem we have $\alpha_{\tau}\neq 0$ and $\beta_{\tau}\neq 0$. Then there exists $\omega_{\tau}\in\mathbb{C}^{*}$ such that
$\Re\big(\omega_{\tau}\alpha_{\tau})\Re\big(\omega_{\tau}\beta_{\tau})>0\quad\text{or}\quad\Im\big(\omega_{\tau}\alpha_{\tau})\Im\big(\omega_{\tau}\beta_{\tau})>0.$
Suppose that
$\Re\big(\omega_{\tau}\alpha_{\tau})\Re\big(\omega_{\tau}\beta_{\tau})>0.$ (9)
The proof is analogous for the imaginary part. Fix $\omega_{\tau}$ satisfying the previous equation and define
$\tilde{\delta}_{1+i\tau}(t):=\Re\big(\omega_{\tau}\phi_{\eta,\tau}(t)\big)\quad\text{and}\quad\tilde{\delta}(t):=\Re\big(\omega_{\tau}\delta_{\eta,s}(t)\big).$ (10)
Equation (8) implies
$\Big|\int_{1}^{t}\,u^{-1}\tilde{\delta}_{1+i\tau}(u)\,du+\Re\big(\omega_{\tau}\alpha_{\tau})t-\Re\big(\omega_{\tau}\beta_{\tau})\,\ln(t)\Big|<M_{\tau}\,|\omega_{\tau}|,\quad\forall t\geq 1,$ (11)
For every $c>1$, denote
$\mathcal{A}_{c}(t):=\frac{1}{2}+\frac{1}{\pi}\arctan(t-c);\quad\forall t\geq 1.$
For every $c>1$ and $g\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$, consider the following transformation
$\Phi_{c}[g](t):=c\,t^{-1}\int_{1}^{t}\mathcal{A}_{c}(u)u^{-1}g(u)\,du,\quad\forall t\geq 1.$
Since $\lim_{t\to+\infty}\mathcal{A}_{c}(t)=1$ for every $c>1$, by equation (11), we have
$\forall c>1:\quad\lim_{t\to+\infty}\Big|t^{-1}\int_{1}^{t}\mathcal{A}_{c}(u)u^{-1}\tilde{\delta}_{1+i\tau}(u)\,du\Big|=\big|\Re\big(\omega_{\tau}\alpha_{\tau})\big|>0.$ (12)
Using equation (9), there exist a sufficiently large $c_{s}>1$ and $t_{c_{s}}>1$ such that for all $c>c_{s}$ and $t>t_{c_{s}}$ we have
$\Big|t^{-1}\int_{1}^{t}\mathcal{A}_{c}(u)u^{-1}\tilde{\delta}_{1+i\tau}(u)\,du\Big|=\max_{z\in[1,t]}\Big|z^{-1}\int_{1}^{z}\mathcal{A}_{c}(u)u^{-1}\tilde{\delta}_{1+i\tau}(u)\,du\Big|,$ (13)
Implies
$\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|=\max_{z\in[1,t]}\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](z)\Big|,\quad\forall c>c_{s},\quad\forall t>t_{c_{s}}.$ (14)
The purpose of introducing $\mathcal{A}_{c}$ is to obtain the previous equality. It ensures that the peaks of the function $t\mapsto\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|$ are bounded above by the limiting value which is equal to $c|\Re\big(\omega_{\tau}\alpha_{\tau})|$.
By the Proposition 7, for all $t\geq 1$ we have
$\frac{t^{-1}}{\sigma(1-\sigma)}\Big(\tilde{\delta}_{1+i\tau}(t)$ $-\tilde{\delta}_{s}(t)\Big)$
$=t^{-\sigma}\int_{1}^{t}\,u^{-2(1-\sigma)}\int_{1}^{u}\,v^{-1-\sigma}\tilde{\delta}_{1+i\tau}(v)\,dv\ du,$
Applying the transformation $\Phi_{c}$,
$\frac{t}{\sigma(1-\sigma)}\Big(\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)-\Phi_{c}[\tilde{\delta}_{s}](t)\Big)$
$=c\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,v^{-2(1-\sigma)}\int_{1}^{v}\,\frac{z^{-\sigma}}{\mathcal{A}_{c}(z)}\mathcal{A}_{c}(z)z^{-1}\tilde{\delta}_{1+i\tau}(z)\,dz\,dv\ du,$
Use the integration by part formula,
$\frac{t}{\sigma(1-\sigma)}\Big(\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)-\Phi_{c}[\tilde{\delta}_{s}](t)\Big)$
$=\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,\frac{v^{-1+\sigma}}{\mathcal{A}_{c}(v)}\Phi_{c}[\tilde{\delta}_{1+i\tau}](v)\,dv\,du$
$-\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,v^{-2(1-\sigma)}\int_{1}^{v}z\Big(\frac{d}{dz}\frac{z^{-\sigma}}{\mathcal{A}_{c}(z)}\Big)\,\Phi_{c}[\tilde{\delta}_{1+i\tau}](z)\,dz\,dv\,du,$
We have
$\frac{4}{\pi}\frac{d}{dz}\frac{z^{-\sigma}}{\mathcal{A}_{c}(z)}=-\frac{\sigma\,z^{-\sigma-1}\left(\frac{\pi}{2}+\arctan(z-c)\right)+\frac{z^{-\sigma}}{1+(z-c)^{2}}}{\left(\frac{\pi}{2}+\arctan(z-c)\right)^{2}}<0,\quad\forall z\geq 1.$
By equation (14), for all $c>c_{s}$ and for all $t\geq t_{c_{s}}$ , we obtain
$t\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|<t\Big|\Phi_{c}[\tilde{\delta}_{s}](t)\Big|+\sigma(1-\sigma)\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|K(\sigma,c,t),$ (15)
where
$K(\sigma,c,t)$ $:=\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,\frac{v^{-1+\sigma}}{\mathcal{A}_{c}(v)}\,dv\,du$
$-\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,v^{-2(1-\sigma)}\int_{1}^{v}z\Big(\frac{d}{dz}\frac{z^{-\sigma}}{\mathcal{A}_{c}(z)}\Big)\,\,dz\,dv\,du,\quad\forall t\geq 1.$
We prove the theorem by contradiction. Suppose that $\mu_{\eta}(s)=\mu_{\eta}(1-\overline{s})=0$. Using the notation of $\delta_{\eta,s}$ given in Notation 6,it follows from the second item of the Lemma 3 that there exists a constant $C_{s}>0$ such that such that
$\left|\delta_{\eta,s}(t)\right|<C_{s},\quad\forall t\geq 1.$
Using the following compact convergence
$\lim_{c\to+\infty}c\,\mathcal{A}_{c}(u)=\frac{1}{\pi},$
We obtain
$\lim_{c\to+\infty}\Big|t\Phi_{c}[{\delta}_{\eta,s}](t)\Big|=\lim_{c\to+\infty}\Big|c\int_{1}^{t}\,\mathcal{A}_{c}(u)u^{-1}\delta_{\eta,s}(u)\,du\Big|$ $\leq\frac{1}{\pi}\,C_{s}\,\ln(t),\quad\forall t\geq 1.$
By the notation of $\tilde{\delta}_{s}$ in equation (10), we get
$\lim_{c\to+\infty}\Big|t\Phi_{c}[\tilde{\delta}_{s}](t)\Big|<\frac{1}{\pi}|\omega_{\tau}|\,C_{s}\,\ln(t),\quad\forall t\geq 1.$
Equation (15) gives
$\lim_{c\to+\infty}\frac{\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|}{\ln(t)}\Big(t-\sigma(1-\sigma)K(\sigma,c,t)\Big)<\frac{1}{\pi}|\omega_{\tau}|\,C_{s},\quad\quad\forall t\geq t_{c_{s}}.$ (16)
Now, let us examine the asymptotic behavior of $K(\sigma,c,t)$. We recall the definition
$\mathcal{A}_{c}(t)=\frac{1}{2}+\frac{1}{\pi}\arctan(t-c),\qquad t\geq 1,$
and
$K(\sigma,c,t)$ $=\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,\frac{v^{-1+\sigma}}{\mathcal{A}_{c}(v)}\,dv\,du$
$-\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,v^{-2(1-\sigma)}\int_{1}^{v}z\Big(\frac{d}{dz}\frac{z^{-\sigma}}{\mathcal{A}_{c}(z)}\Big)\,\,dz\,dv\,du,\quad\forall t\geq 1.$
Use the integration by parts formula to obtain
$K(\sigma,c,t)=\frac{1}{2\sigma-1}\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}W(u)\,du,$ (17)
with
$W(u):=\dfrac{1}{\mathcal{A}_{c}(1)}(u^{-1+2\sigma}-1)+u^{-1+2\sigma}\int_{1}^{u}\dfrac{v^{-\sigma}}{\mathcal{A}_{c}(v)}\,dv-\int_{1}^{u}\frac{v^{-1+\sigma}}{\mathcal{A}_{c}(v)}\,dv.$
Use the integration by part formula, to get
$\mathcal{A}_{c}(u)u^{-\sigma}W(u)$ $=\frac{1}{1-\sigma}-\frac{1}{\sigma}$
$-\frac{\mathcal{A}_{c}(u)}{1-\sigma}u^{-1+\sigma}\Big[\dfrac{\sigma}{\mathcal{A}_{c}(1)}+\int_{1}^{u}v^{1-\sigma}\frac{d}{dv}\dfrac{1}{\mathcal{A}_{c}(v)}\,dv\Big]$
$+\frac{\mathcal{A}_{c}(u)}{\sigma}u^{-\sigma}\Big[\dfrac{1-\sigma}{\mathcal{A}_{c}(1)}+\int_{1}^{u}v^{\sigma}\frac{d}{dv}\dfrac{1}{\mathcal{A}_{c}(v)}\,dv\Big]$
Since
$\frac{d}{du}\mathcal{A}_{c}(u)=\frac{1}{\pi}\frac{1}{1+(u-c)^{2}},\quad\forall u\geq 1,$
It follows that
$\mathcal{A}_{c}(u)u^{-\sigma}W(u)$ $=\frac{1}{1-\sigma}-\frac{1}{\sigma}$
$-\frac{\mathcal{A}_{c}(u)}{1-\sigma}u^{-1+\sigma}\Big[\dfrac{\sigma}{\mathcal{A}_{c}(1)}-\frac{1}{\pi}\int_{1}^{u}\dfrac{1}{(\mathcal{A}_{c}(v))^{2}}\frac{v^{1-\sigma}}{1+(v-c)^{2}}\,dv\Big]$
$+\frac{\mathcal{A}_{c}(u)}{\sigma}u^{-\sigma}\Big[\dfrac{1-\sigma}{\mathcal{A}_{c}(1)}-\frac{1}{\pi}\int_{1}^{u}\dfrac{1}{(\mathcal{A}_{c}(v))^{2}}\frac{v^{\sigma}}{1+(v-c)^{2}}\,dv\Big]$
Then
$\lim_{c\to+\infty}\mathcal{A}_{c}(u)u^{-\sigma}W(u)=\frac{2\sigma-1}{\sigma(1-\sigma)}-\frac{\sigma}{1-\sigma}u^{-1+\sigma}+\frac{1-\sigma}{\sigma}u^{-\sigma}$
Inserting these into the integral for $K(\sigma,c,t)$ given by equation (17), we obtain the following equation
$\lim_{c\to+\infty}\Big(t-\sigma(1-\sigma)K(\sigma,c,t)\Big)=\frac{1}{2\sigma-1}\,\Big(\sigma t^{\sigma}-(1-\sigma)t^{1-\sigma}\Big).$
Using the following compact convergence
$\lim_{c\to+\infty}c\,\mathcal{A}_{c}(u)=\frac{1}{\pi},$
Thanks to equation (11), we get
$\lim_{t\to+\infty}\lim_{c\to+\infty}\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|=\frac{1}{\pi}\big|\Re\big(\omega_{\tau}\alpha_{\tau})\big|>0.$
Since $\sigma\in(0,1)\setminus\frac{1}{2}$, we obtain
$\lim_{t\to+\infty}\lim_{c\to+\infty}\frac{\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|}{\ln(t)}\Big(t-\sigma(1-\sigma)K(\sigma,c,t)\Big)=+\infty.$
We obtain a contradiction with equation (16). ∎